OpenStudy (anonymous):

xy^2+4xy=10 find dy/dx

OpenStudy (bahrom7893):

x*2yy'+y^2 + 4xy' + 4y = 0 x*2yy'+4xy' = -(4y+y^2) y'(2xy+4x)=-(4y+y^2)

OpenStudy (bahrom7893):

y'=dy/dx = -y(4+y)/[2x(y+2)]

OpenStudy (bahrom7893):

check the arithmetic

OpenStudy (anonymous):

I'm a little confused what are the * for?

OpenStudy (anonymous):

darn it im too slow :(\[\frac{d}{dx} \left(xy^{2}\right) +\frac{d}{dx} \left(4xy\right) =\frac{d}{dx} \left(10\right) => x2yy' +y^2 + 4xy'+y =0\]

OpenStudy (anonymous):

you use the product rule

OpenStudy (bahrom7893):

* means times

OpenStudy (anonymous):

I'm wretched at caluculus. Thank you so much